OpenStudy (ttop0816):

easy question! Which of the following lines has a slope of -? x + 2y = 0 x - 2y = 0 -x + 2y = 0

OpenStudy (anonymous):

Do you mean you want to know which of these have a negative slope?

OpenStudy (ttop0816):

slope of -1/2!

OpenStudy (anonymous):

So you can learn how to do it, what is it you exactly don't understand? :)

OpenStudy (anonymous):

Do you know slope-intercept form and standard form?

hartnn (hartnn):

try to bring every equation in the form of y=mx by isolating the y then m is the slope

OpenStudy (ttop0816):

x + 2y = 0?

hartnn (hartnn):

first subtract x from both sides, what u get ?

OpenStudy (ttop0816):

i tried to solve the question and i got x + 2y = 0!

hartnn (hartnn):

you mean you got slope of -1/2 for x+2y=0 ??

hartnn (hartnn):

absolutely! :)

OpenStudy (ttop0816):

@hartnn thank you so much!!

hartnn (hartnn):

@ttop0816 welcome so much!!